Jump to main content.


Drake Chemical

EPA Daily Situation Reports for January 1999

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Sunday, January 31, 1999

FOR DATES: From 7:00 a.m. Saturday, Jan. 30, 1999 to 7:00 a.m. Sunday, Jan. 31, 1999.

WEATHER: Mostly sunny. The temperature range was 17 to 38oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 848.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.67 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, lime mixing, ash handling, and ash backfilling. Excavation activities were continued in the corner of the exclusion zone where the ash pad sheet steel meets the utility corridor sheet piling. Excavation was also performed along the exclusion zone sheet piling in preparation for whaler installation. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the whaler shoring system for the sheet piling in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, January 30, 1999

FOR DATES: From 7:00 a.m. Friday, Jan. 29, 1999 to 7:00 a.m. Saturday, Jan. 30, 1999.

WEATHER: Partly Sunny. The temperature range was 28oF to 38oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 833.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, equipment decontamination, debris handling, debris separation, lime mixing, ash handling, and ash backfilling. Excavation activities were continued along the southern face of the exclusion zone in the vicinity of the former debris separation building. Excavation was also performed along the exclusion zone sheet piling in preparation for whaler installation. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the whaler shoring system for the sheet piling in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, January 29, 1999

FOR DATES: From 7:00 a.m. Thursday, Jan. 28, 1999 to 7:00 a.m. Friday, Jan. 29, 1999.

WEATHER: Sunny. The temperature range was 30oF to 49oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 780.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 32.7 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, equipment decontamination, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were continued adjacent to the utility corridor sheetpiling. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-01-23-99-342) for the contents of bin 1. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical report (fly-ash-01-24-99-068) for the contents of bin B. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of cadmium, chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the whaler shoring system for the sheet piling in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, January 28, 1999

FOR DATES: From 7:00 a.m. Wednesday, Jan. 27, 1999 to 7:00 a.m. Thursday, Jan. 28, 1999.

WEATHER: Partly cloudy. The temperature range was 24oF to 36oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 762.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were performed adjacent to the ash stabilization building. Nestlerode Construction continued to install whalers on the sheet piling adjacent to the WWTP. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-01-23-99-342, bot-ash-01-18-99-338 and bot-ash-01-20-99-339) for the contents of bin 2, bin 7 and bin 6, respectively. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the whaler shoring system for the sheet piling in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Kurt Davis (AIR).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, January 27, 1999

FOR DATES: From 7:00 a.m. Tuesday, Jan. 26, 1999 to 7:00 a.m. Wednesday, Jan. 27, 1999.

WEATHER: Partly cloudy. The temperature range was 22oF to 40oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 752.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31.5 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, WWTP weekly NPDES sampling, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were performed near the ash stabilization building in order to uncover site sheet piling. Nestlerode Construction began installing whalers on the sheet piling adjacent to the WWTP. Continued with the real time perimeter air monitoring (PAM) program and collected all of the air sampling filters from the OSHA method 93 weekly sampling event for BNA.

Received bottom ash analytical report (bot-ash-01-22-99-341) for the contents of bin 3. Bottom ash analytical test results indicated that the thermally treated soil did not meet the project requirements for backfilling on-site. Test results indicated that the concentration of B-Naphthylamine (BNA) was slightly higher than the allowable organic criteria concentration for BNA, as a result the contents of bin 3 were placed back in the exclusion zone for further thermal treatment.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the whaler shoring system for the sheet piling in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs (PADEP), Jared Olsen (USACE), Glenn Phelps (OHM/IT), Kurt Davis (AIR).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Tuesday, January 26, 1999

FOR DATES: From 7:00 a.m. Monday, Jan. 25, 1999 to 7:00 a.m. Tuesday, Jan. 26, 1999.

WEATHER: Overcast with periods of rain. The temperature range was 31oF to 38oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.6 hours and 769.0 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is primarily attributed to the performance of a full calibration of the weighbelt. The TDF processed contaminated soil at an average rate of 32.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, drum handling, lime mixing, ash handling, and ash backfilling. Excavation activities were continued in the general vicinity of the former debris separation building. Nestlerode Construction finished installing H-piles along the sheet steel adjacent to sediment basin 2. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-01-21-99-340) for the contents of bin 5. Bottom ash analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the whaler shoring system for the sheet piling in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Monday, January 25, 1999

FOR DATES: From 7:00 a.m. Sunday, Jan. 24, 1999 to 7:00 a.m. Monday, Jan. 25, 1999.

WEATHER: Overcast with rain. The temperature range was 36oF to 46oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 807.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, excavation, debris handling, debris separation, lime mixing, ash handling, & ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the H-pile shoring system for the sheet steel in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for William Werntges (USACE)


Sunday, January 24, 1999

FOR DATES: From 7:00 a.m. Saturday, Jan. 23, 1999 to 7:00 a.m. Sunday, Jan. 24, 1999.

WEATHER: Overcast with rain. The temperature range was 36oF to 44oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 771.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 32.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, excavation, debris handling, debris separation, lime mixing, ash handling, ash backfilling, and waste water treatment. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the H-pile shoring system for the sheet steel in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, January 23, 1999

FOR DATES: From 7:00 a.m. Friday, Jan. 22, 1999 to 7:00 a.m. Saturday, Jan. 23, 1999.

WEATHER: Overcast with rain. The temperature range was 36oF to 41oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 793.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, lime mixing, ash handling, and ash backfilling. The top of the sheet steel adjacent to sediment basin 2 was uncovered on this date in order to prepare the area for installation of the H-pile shoring system. Nestlerode Construction continued to drive H-piles adjacent to the ash stabilization building sheet steel. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the H-pile shoring system for the sheet steel in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, January 22, 1999

FOR DATES: From 7:00 a.m. Thursday, Jan. 21, 1999 to 7:00 a.m. Friday, Jan. 22, 1999.

WEATHER: Overcast and rainy. The temperature range was 26oF to 36oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 23.9 hours and 810.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.9 tons per hour while on-line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, wood shredding, lime mixing, debris separation, debris handling, rock cleaning, rock backfilling, ash handling and ash backfilling. No new excavation was performed on this date. The contractor continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-01-16-99-336 and bot-ash-01-17-99-337) for the contents of bin 2 and bin 1, respectively. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site. Received fly ash analytical report (fly-ash-01-15-99-067) for the contents of bin A. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentration of cadmium, chromium and selenium being slightly higher than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, shoring system installation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, January 21, 1999

FOR DATES: From 7:00 a.m. Wednesday, Jan. 20, 1999 to 7:00 a.m. Thursday, Jan. 21, 1999.

WEATHER: Clear and mild. The temperature range was 24oF to 45oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 22.0 hours and 712.4 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is primarily attributed to the stoppage of feed in order to replace a roller on the weighbelt conveyor system. The TDF processed contaminated soil at an average rate of 32.4 tons per hour while on-line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, wood shredding, lime mixing, debris separation, debris handling, rock cleaning, rock backfilling, ash handling and ash backfilling. Excavation activities continued on the eastern side of the exclusion zone adjacent to the Conrail sheet piling. The contractor continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, shoring system installation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandra Downs (PADEP)

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, January 20, 1999

FOR DATES: From 7:00 a.m. Tuesday, Jan. 19, 1999 to 7:00 a.m. Wednesday, Jan. 20, 1999.

WEATHER: Partly cloudy and windy. The temperature range was 30oF to 44oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 18.0 hours and 519.3 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is attributed to maintenance being performed on the Wet Ash Drag conveyor. The TDF processed contaminated soil at an average rate of 28.9 tons per hour while on-line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance, hazardous waste feed to the incinerator resumed at 1253 hours on January 19, 1999. The contractor continued Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris separation, debris handling, ash handling and ash backfilling. The contractor sampled the contents of 9 potential drums for fingerprint analyses. The drums were then overpacked and placed in a temporary storage area in the exclusion zone awaiting the results of the fingerprint analyses. Excavation activities continued on the eastern side of the exclusion zone adjacent to the Conrail sheet piling. The contractor continued with the real time perimeter air monitoring (PAM) program and began the weekly OHSA Method 93 24-hour air sampling event at the on-site and off-site air monitoring stations. Nestlerode Construction continued to drive H-piles at the sheet steel in the exclusion zone.

Received bottom ash analytical reports (bot-ash-01-13-99-333A) for the contents of bin 5, (bot-ash-01-14-99-334) for the contents of bin 4 and (bot-ash-01-15-99-335) for the contents of bin 3. Analytical test results indicated that the thermally treated soil sampled from bin 4 met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Analytical test results for the contents of bin 3 and bin 5 indicated that thermally treated soil met the project requirements for restricted backfilling on-site (i.e above elevation 553) due to the concentrations of arsenic and lead in bin 3, and arsenic in bin 5 being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, shoring system installation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs and Larry Newcomer (PADEP), Dr. Richard Stehouwer (Penn State University).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Tuesday, January 19, 1999

FOR DATES: From 7:00 a.m. Monday, Jan. 18, 1999 to 7:00 a.m. Tuesday, Jan. 19, 1999.

WEATHER: Overcast and windy with rain. The temperature range was 36oF to 40oF.

I. CURRENT SITUATION: Feed to the Thermal Destruction Facility (TDF) is currently off-line while preventative maintenance is being performed on the Dry Ash Drag (DAD) and Wet Ash Drag (WAD) conveyors. OHM continues to work 24 hours a day, seven days a week maintaining the TDF, which is currently in a hot hold mode. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance. The TDF preventative maintenance included repairs to the double dump valves, repairs to the DAD solenoids and idler bearings, and repairs to the WAD drive chain and the replacement of several bent flights. The contractor continued Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris separation, debris handling, ash handling and ash backfilling. Excavation activities were continued on the eastern side of the exclusion zone adjacent to the Conrail sheet piling. The contractor continued with the real time perimeter air monitoring (PAM) program. Nestlerode Construction began driving H-piles along the sheet steel in the exclusion zone.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, shoring system installation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: Repairs to the WAD were required when the drive chain broke during the preparation for restart of the TDF. Also, during inspection of the WAD drive chain, several flights were observed to be bent and will be replaced. Hazardous waste feed to the incinerator is scheduled to resume in the early afternoon, once repairs to the WAD are complete and the kiln reaches operating temperatures.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Monday, January 18, 1999

FOR DATES: From 7:00 a.m. Sunday, Jan. 17, 1999 to 7:00 a.m. Monday, Jan. 18, 1999.

WEATHER: Mostly sunny and warm. The temperature range was 23oF to 43oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 19.3 hours and 553.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 28.7 tons per hour while on line. The plant is presently in a hot hold condition due in order to allow repairs to the double dump valve. Repairs are presently being made and resumption of feed is expected by noon Monday. Related ongoing site activities feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, lime mixing, snow removal, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, installing the sheetsteel shoring system, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: Repair of double dump valve should be completed by noon Monday. Maintenance, feed screw measurements and down time repairs is being made while plant is in this hot hold condition.

V. NOTED VISITORS: None. Down time notifications were made to Gregg Crystall, Sandy Downs and George Drumbor.

PREPARED BY: John Ricard for William Werntges (USACE)


Sunday, January 17, 1999

FOR DATES: From 7:00 a.m. Saturday, Jan. 16, 1999 to 7:00 a.m. Sunday, Jan. 17, 1999.

WEATHER: Mostly sunny and breezy. The temperature range was 18oF to 43oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 677.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 28.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, lime mixing, snow removal, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-1-11-99-331) for bin 7 and (bot-ash-1-12-99-332) for bin 6. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547.0).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, installing the sheetsteel shoring system, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, January 16, 1999

FOR DATES: From 7:00 a.m. Friday, Jan. 15, 1999 to 7:00 a.m. Saturday, Jan. 16, 1999.

WEATHER: Snow and freezing rain; clearing later. The temperature range was 17oF to 27oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 644.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 27.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, lime mixing, snow removal, ash handling, and ash backfilling. Excavation activities were continued along the southern face of the exclusion zone in the general vicinity of the debris separation building; no drums were encountered on this date. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, installing the sheetsteel shoring system, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, January 15, 19999

FOR DATES: From 7:00 a.m. Thursday, Jan. 14, 1999 to 7:00 a.m. Friday, Jan. 15, 1999.

WEATHER: Freezing rain and snow, heavy at times. The temperature range was 9oF to 20oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 746.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris separation, debris handling, drum handling, rock cleaning, rock backfilling, ash handling and ash backfilling. The contractor continued excavation activities adjacent to the corridor sheetpiling; no drums were encountered. The contractor continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-01-10-99-330) for the contents of bin 1. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical report (fly-ash-01-10-99-066) for the contents of bin E. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of arsenic, chromium and selenium being greater than the PADEP drinking water standard but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, shoring system installation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: As a result of the winter storm which passed through the region, OHM has indicated that the liquid oxygen deliveries may be delayed due to impassable roads. If this occurs, waste feed to the TDF may be suspended and the unit will go into a hot hold mode until deliveries resume.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, Janury 14, 1999

FOR DATES: From 7:00 a.m. Wednesday, Jan. 13, 1999 to 7:00 a.m. Thursday, Jan. 14, 1999.

WEATHER: Overcast with freezing rain and snow. The temperature range was 5oF to 32oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 835.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris separation, debris handling, drum handling, rock cleaning, rock backfilling, ash handling and ash backfilling. The contractor continued excavation activities adjacent to the corridor sheetpiles; several potential drums, drum carcasses were unearthed along with demolition debris. The contractor continued with the real time perimeter air monitoring (PAM) program and initiated the OHSA Method 93 24-hour air sampling event for BNA at the on-site and off-site air monitoring stations. Received bottom ash analytical reports (bot-ash-01-07-99-327, bot-ash-01-08-99-328 and bot-ash-01-09-99-329) for bin 4, bin 3 and bin 2, respectively. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, shoring system installation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: PADEP was notified that the steam scrubber was off-line from approximately 1100 hours to 1255 hours to replace a faulty valve on the re-circulation system.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, January 13, 1999

FOR DATES: From 7:00 a.m. Tuesday, Jan. 12, 1999 to 7:00 a.m. Wednesday, Jan. 13, 1999.

WEATHER: Overcast with freezing rain. The temperature range was 22oF to 32oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 18.3 hours and 603.8 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is attributed to feed stoppage during the monthly interlock checks. Once the interlock tests had been completed at 1100 hours, the TDF remained in "hot hold" while thermal crews performed preventative maintenance to the dual feed screws. The TDF processed contaminated soil at an average rate of 33.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance, monthly TDF interlock checks, monthly TDF equipment calibrations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris separation, debris handling, rock handling, ash handling and ash backfilling. The remainder of the recently excavated intact drums were overpacked and sampled to determine chemical compatibility and reactivity. The contractor continued with the real time perimeter air monitoring (PAM) program. The contractor continued excavation activities adjacent to the corridor sheetpiling. Nestlerode Construction mobilized to the exclusion zone in order to begin installing the sheetsteel shoring system.

Hazardous waste feed to the incinerator was stopped at 0834 hours on January 12, 1999 in order to perform the monthly interlock checks, which test the incinerator's automatic shut down systems. As part of the interlock checks, a planned Thermal Relief Valve (TRV) opening was initiated at 1058 hours. Prior to the TRV opening, hazardous waste feed to the incinerator had been off for approximately 2.5 hours and the proper regulatory authorities were notified.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs, Larry Newcomer and Mike Welch (PADEP).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Tuesday, January 12, 1999

FOR DATES: From 7:00 a.m. Monday, Jan. 11, 1999 to 7:00 a.m. Tuesday, Jan. 12, 1999.

WEATHER: Partly cloudy with snow in the evening. The temperature range was 11oF to 22oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 689.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 29.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation. Nestlerode Construction is on-site and will be installing the H-pile shoring system for the sheet steel in the exclusion zone.

Hazardous waste feed will be stopped at 0830 hours on January 12, 1999 in order to perform monthly interlock checks on the incinerators automatic shut down systems. As part of the interlock checks, a planned Thermal Relief Valve (TRV) opening will be initiated during the later morning hours. Prior to the TRV opening, the kiln will be purged of all hazardous waste feed and the proper regulatory authorities will be notified.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris separation, debris handling, rock cleaning, rock backfilling, ash handling and ash backfilling. Approximately two feet of shot rock was removed from areas along the sheet steel in the exclusion zone in preparation for pile driving to be performed on Tuesday, January 12, 1999. Excavation was performed on the southern embankment face of the exclusion zone. The contractor continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-1-6-99-236) for bin 5. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical report (fly-ash-1-4-99-065) for bin D. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. greater than elevation 553) due to the concentrations of cadmium and chromium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Monday, January 11, 1999

FOR DATES: From 7:00 a.m. Sunday, Jan. 10, 1999 to 7:00 a.m. Monday, Jan. 11, 1999.

WEATHER: Partly cloudy. The temperature range was 6oF to 22oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 752.0 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris separation, debris handling, rock backfilling, ash handling and ash backfilling. The contractor continued to construct a heated enclosure for the decontamination pad. Approximately two feet of shot rock was removed from areas along the sheet steel in the exclusion zone in preparation for pile driving to be performed on Tuesday, January 12, 1999. The contractor continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. The contractor will be performing monthly interlock checks and calibrations on Tuesday, January 12, 1999, this will include a planned opening of the thermal relief vent.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Sunday, January 19, 1999

FOR DATES: From 7:00 a.m. Saturday, Jan. 9, 1999 to 7:00 a.m. Sunday, Jan. 10, 1999.

WEATHER: Snow, sleet, rain, and freezing rain. The temperature range was 8oF to 33oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 785.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 32.8 tons per hour while on line. Down time was contributed to the daily zero weigh belt calibration. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, ash backfilling and site housekeeping. The contractor continued to erect a heated enclosure for the decontamination pad. Excavation activities were continued along the southern face of the exclusion zone in the general vicinity of the former debris separation building location. Excavation was also performed along the ash pad sheet steel in order to remove the large shot rock at the proposed H-pile locations. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-01-05-99-325) for bin 6. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547.0).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the H-pile shoring system for the sheet steel in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, January 9, 1999

FOR DATES: From 7:00 a.m. Friday, Jan. 8, 1999 to 7:00 a.m. Saturday, Jan. 9, 1999.

WEATHER: Snow, sleet and freezing rain. The temperature range was 10oF to 30oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 759.0 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 32.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, ash backfilling and site housekeeping. The contractor continued to erect a heated enclosure for the decontamination pad. Excavation activities were continued along the southern face of the exclusion zone in the general vicinity of the former debris separation building location. Excavation was also performed along the ash pad sheet steel in order to remove the large shot rock at the proposed H-pile locations. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-01-04-99-324) for bin 7. Analytical test results indicated that the thermally treated soil met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentration of lead being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the H-pile shoring system for the sheet steel in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, January 8, 1999

FOR DATES: From 7:00 a.m. Thursday, Jan. 7, 1999 to 7:00 a.m. Friday, Jan. 8, 1999.

WEATHER: Partly cloudy. The temperature range was 5oF to 24oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 668.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 27.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris separation, debris handling, ash handling and ash backfilling. The contractor continued to construct a heated enclosure for the decontamination pad. Excavation continued on the southern end of the exclusion zone in the vicinity of the former debris separation building; no intact drums were encountered. The rockwash pad and approximately two feet of shot rock were removed from areas along the sheet steel in the exclusion zone in preparation for pile driving to be performed on Monday, January 11, 1998. The contractor continued with the real time perimeter air monitoring (PAM) program and the weekly OSHA method 93 24-hour sampling event was initiated at the on-site and off-site air monitoring stations.

Received bottom ash analytical report (bot-ash-01-03-99-323) for bin 1. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical report (fly-ash-12-30-98-064) for bin C. Test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install an H-pile shoring system for the sheet steel located in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Christopher Cain (Montgomery Watson)

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, January 7, 1999

FOR DATES: From 7:00 a.m. Wednesday, Jan. 6, 1999 to 7:00 a.m. Thursday, Jan. 7, 1999.

WEATHER: Overcast with snow in the evening. The temperature range was 3oF to 25oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 734.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 30.7 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris separation, debris handling, wood shredding, rock cleaning, rock backfilling, ash handling and ash backfilling. Excavation continued on the southern end of the exclusion zone in the vicinity of the former debris separation building. Several drum carcasses and several intact drums were unearthed during excavation activities. The intact drums were separated for sampling and overpacking. Two rolloff containers containing non-hazardous metal debris and PPE were shipped off-site to a hazardous waste landfill. The contractor continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-01-02-99-322) for bin 2. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Contracting will be on-site to install a shoring system for the sheet steel located in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs (PADEP).

PREPARED BY: Steve Gibson for William Werntges(USACE)


Wednesday, January 6, 1999

FOR DATES: From 7:00 a.m. Tuesday, Jan. 5, 1999 to 7:00 a.m. Wednesday, Jan. 6, 1999.

WEATHER: Clear. The temperature range was 3oF to 20oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 703.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 29.7 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris separation, debris handling, wood shredding, rock cleaning, rock backfilling, ash handling and ash backfilling. Excavation continued on the southern end of the exclusion zone adjacent to the utility corridor. Several drum carcasses and several intact drums were unearthed during excavation activities. The intact drums were separated for sampling and overpacking. The contractor continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-01-01-99-321) for bin 3. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs and Larry Newcomer (PADEP).

PREPARED BY: Steve Gibson for William Werntges(USACE)


Tuesday, January 5, 1999

FOR DATES: From 7:00 a.m. Monday, Jan. 4, 1999 to 7:00 a.m. Tuesday, Jan. 5, 1999.

WEATHER: Partly cloudy. The temperature range was 10oF to 20oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 710.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 30.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris separation, debris handling, wood shredding, ash handling and ash backfilling. No new excavation was performed on this date. Continued with the real time perimeter air monitoring (PAM) program. Contractor continues to mix both hydrated lime and quick lime to site soils to improve material handling properties. Two roll-offs of PPE and miscellaneous metal debris was shipped off-site to a hazardous waste landfill.

Received bottom ash analytical reports (bot-ash-12-29-98-318, bot-ash-12-30-98-319 and bot-ash-12-31-98-320) for bin 6, bin 5 and bin 4, respectively. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges(USACE)


Monday, January 4, 1999

FOR DATES: From 7:00 a.m. Sunday, Jan. 3, 1999 to 7:00 a.m. Monday, Jan. 4, 1999.

WEATHER: Snow, sleet, freezing rain am, cold clear pm. The temperature range was 18oF to 35oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.3 hours and 748.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 32.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris separation, debris handling, ash handling and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program. No new ash analytical reports received during this period. Downtime during this period attributed to daily weigh belt calibration and the flame out of an SCC burner, both of which stopped feed to the incinerator briefly. Contractor performed snow and ice removal during periods of inclement weather. Contractor continues to mix both hydrated lime and quick lime to site soils to improve material handling properties.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Chad Thompson for William Werntges(USACE)


Sunday, January 3, 1999

FOR DATES: From 7:00 a.m. Saturday, Jan. 2, 1999 to 7:00 a.m. Sunday, Jan. 3, 1999.

WEATHER: A mixture of snow, sleet, freezing rain, and rain. The temperature range was 6oF to 29oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 743.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31.2 tons per hour while on line. Down time was attributed to the daily zero weigh belt calibration. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling and separation, mixing lime, ash handling, and ash back filling. Minor excavation and grading was performed in the vicinity of the former debris separation building. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-12-28-98-317). Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted back filling on site.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None.

PREPARED BY: Richard L. Lupold for Chad Thompson (USACE)


Saturday, January 2, 1999

FOR DATES: From 7:00 a.m. Friday, Jan. 1, 1999 to 7:00 a.m. Saturday, Jan. 2, 1999.

WEATHER: Partly cloudy with light snow. The temperature range was 5oF to 24oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 836.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, rock cleaning, rock backfilling and ash handling. Minor excavation and grading was performed in the vicinity of the former debris separation building. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlrode Contracting will be on-site later next week in order to install a shoring system for the exclusion zone sheet steel.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Friday, January 1, 1999

FOR DATES: From 7:00 a.m. Thursday, Dec. 31, 1998 to 7:00 a.m. Friday, Jan. 1, 1999.

WEATHER: Windy, snow showers. The temperature range was 16oF to 23oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours 55 minutes and 846.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, drum handling, excavation, feed preparation, lime mixing, debris separation, debris handling, wood shredding, ash handling and ash backfilling. Continued to sample and overpack drums. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical reports (bot-ash-12-26-98-316). Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Chad Thompson (USACE)

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.